# Long-Term Mapping of the Douro River Plume with Multi-Agent Reinforcement Learning

### v2

@@ II Problem Formulation
where $\Omega\subset\mathbb{R}^{2}$ denotes a planar polygonal domain and $\mathbb{R}_{+}$ is time. In this work, we consider 2-dimensional mapping of the plume, which is usually considered the region in $\Omega$ where salinity is lower than the open-ocean background level $f_{ocn}$ ($\approx 35$ psu). For a chosen threshold $\zeta>0$, we can define the plume at time $t$ as
@@
Our objective in this work is to construct estimates $\hat{f}$ of the salinity field $f$ over time, based on measurements collected by a fleet of NN AUVs deployed over $\Omega$. A vehicle $n$ follows the following differential equation mobility model [24]:
@@
To construct the estimate $\hat{f}$ of the unknown function $f$ over space and time, we want to collect measurements of the function $f$ with the AUVs. Given that the AUVs move according to (3), the measurements that we can use for the estimation are constrained in space and time by their NN trajectories. To formally state our problem formulation, we need to introduce some notation. Let us first define a vector of discrete time slots $\sigma=[0,\Delta,2\Delta,...,T\Delta]$ which we index with $\mathcal{K}=\{0,1,...,T\}$. In the rest of the paper, we are interested in updating the estimate $\hat{f}$ over the slots $\sigma[1],...,\sigma[T]$ with a slot granularity of $\Delta=30$ minutes.
@@
where $\epsilon_{x,t}$ is a generic noise term. Given space-time coordinates AA, let $f(A)=\{f(x,t)+\epsilon_{x,t}:(x,t)\in A\}$ be the corresponding measurements, and let $\mathcal{M}_{k}^{N}=\{D_{k}^{N},f_{k}^{N}\}$.
@@ III Plume Monitoring with MARL
In this section, we describe the system architecture and the algorithm we design to solve problem (7). In Fig. 5 we provide a graphical illustration of the system architecture, and Algorithm 1 outlines the main control pipeline. The key rationales behind the system architecture are (i) offloading computation as much as possible from the AUVs to the server, prioritizing their endurance and (ii) relying only on intermittent and extremely low overhead communication. The pipeline at a given time slot $k$ is as follows: first, all agents dwell on the sea surface and establish a connection with the cloud server via Wi-Fi or GSM. During this communication, each AUV n=1,,Nn=1,...,N uplinks its latest measurements $\{Z_{k}^{n},f(Z_{k}^{n})\}$ collected over the past 30 minutes. In our implementation, the number of per-agent per-slot measurements $z_{k,n}$ depends on speed and ocean currents, averaging 55 and limited to a maximum of $10$. Note that transmitting 55 space-time coordinates and the corresponding measurements only involves transmitting 2020 float numbers, which - in the worst case - only requires $160$ bytes (always <500<500 bytes even with wireless transmission overhead). Upon receiving the new measurements $\{Z_{k}^{n},f(Z_{k}^{n})\}_{n=1}^{N}$, the server aggregates them with the previous set, resulting in $\mathcal{M}_{k}^{N}=\{D_{k}^{N},f_{k}^{N}\}=\mathcal{M}_{k-1}^{N}\cup\{Z_{k}^{(n)},f(Z_{k}^{(n)})\}_{n=1}^{N}$ and updates the salinity map estimate $\hat{f}(\cdot,\sigma[k]|\mathcal{M}_{k}^{N})$. Based on $\hat{f}$, the current agent trajectories $D_{k}^{N}$, and a vector containing the wind speed and direction $w_{k}$, the server determines the control policy $\pi(\hat{f},\mathcal{M}_{k}^{N},w_{k})$ for all AUVs n=1,,Nn=1,...,N for the next 30 minutes time slot, which results in direction $\{b_{k}^{(n)}\}_{n=1}^{N}$ and speed levels $\{a_{k}^{(n)}\}_{n=1}^{N}$. At that point, the server downlinks the commands to the AUVs, which resubmerge and resume sampling, starting to build $\{Z_{k+1}^{n},f(Z_{k+1}^{n})\}_{n=1}^{N}$. The estimator $\hat{f}(\cdot,\sigma[k]|\mathcal{M}_{k}^{N})$ and the policy $\pi(\hat{f},\mathcal{M}_{k}^{N},w_{k})$ form the core of our framework. We detail these two components in the following subsections, respectively.
@@ III-B Decision Making Module
To exploit the structure of the action space, we decompose $Q_{\theta}$ into two components, which share a state-value function $V_{\bar{\theta}}(s)$ and two advantage functions $A_{\theta_{1},\text{dir}}(s,b)$, $A_{\theta_{2},\text{spd}}(s,v)$,
@@
where note that $\theta=\{\bar{\theta},\theta_{1},\theta_{2}\}$. Action selection is then:
@@
which lets the agent decide where to go and how fast.
State construction. The state representation $s_{k}^{(n)}$ comprises three components. Following prior work on image-based policy inputs [30], we render both spatial estimates and trajectory traces into a fixed-resolution 33-channels image, compressed by a convolutional neural network (CNN) into a compact embedding. Specifically, the three channels for the $n$-th agent are (i) a compressed image of the GP estimate of the salinity field $\hat{f}(\cdot,\sigma[k])$, (ii) one image marking in white over a black background recent measurement locations of the agent, $Z_{k}^{(n)},Z_{k-1}^{(n)},$... (iii) one image marking in white over a black background recent measurement locations for agent $n$ teammates, $\{Z_{k}^{(l)},Z_{k-1}^{(l)},...\}_{l\neq n}$. The trajectory marks intensities are log-scale weighted by recency. In addition, a wind vector $w_{k}=[\text{angle},\text{speed}]$, processed by a NN to produce an embedding, is concatenated with the rest of the network. See Fig. 5 for an illustration of the state construction.
@@ IV Simulation Study and Results
Dataset. We train our algorithm over a selection of 44 months data from the year 2018 (February, April, October and December) and evaluate over different months from 2016-2018 (see Table I). Overall, we use a total of $\approx 6k$ time frames for the training, with a $30$ minutes time resolution. Each frame spans $\approx 5\times 10^{4}$ grid locations and measurement noise is i.i.d. $\mathcal{N}(0,\,0.01)$.
@@
Multi-agent performance. We now evaluate performance in the multi-agent setting, focusing on: (i) MSE gains provided by the multi-agent coordination when increasing NN, and on the comparison with existing literature benchmarks; (ii) the energy efficiency improvements induced by varying the $\eta_{2}$ reward hyperparameter in (18) and the corresponding MSE-energy tradeoff. As a multi-agent benchmark, we implement the recent work in [32] that proposed an algorithm based on adaptive Voronoi partitioning of the space of interest to track a time-varying process. We empirically tune the exploration-exploitation hyperparameters in [32] and show the best results in terms of MSE. In Fig. 10, we compare the different schemes for $N=3$ and $N=6$ for March 2018. We first note how the adaptive Voronoi partition solution [32] significantly outperforms the strategic rotations solution. We also note that our algorithm provides maps with half of the MSE compared to [32], for both $N=3$ and $N=6$. In Fig. 12, we show a sequence of frames to illustrate the multi-agent control performance of the agents with respect to the salinity field $f$ evolution over a $15$ hour time interval. In Fig. 11 we show the endurance–MSE trade-off obtained by varying the reward weight $\eta_{2}$ in (18), using October 2017 and October 2018 as exemplars. Increasing $\eta_{2}$ biases the policy toward lower propulsion use, thus extending endurance at a modest cost in accuracy, while scaling from $N=3$ to $N=6$ both reduces MSE and more than doubles endurance. Table I summarizes results obtained with $\eta_{2}=50$, and further confirms that our algorithm generalizes across unseen (during training) years and seasonal regimes.

### v3

@@ II Problem Formulation
where $\Omega\subset\mathbb{R}^{2}$ denotes a planar domain and $\mathbb{R}_{+}$ is time. In this work, we consider 2-dimensional mapping of the plume, which is usually considered the region in $\Omega$ where salinity is lower than the open-ocean background level $f_{ocn}$ ($\approx 35$ psu). For a chosen threshold $\zeta>0$, we can define the plume at time $t$ as
@@
Our objective in this work is to construct estimates $\hat{f}$ of the salinity field $f$ over time, based on measurements collected by a fleet of $N$ AUVs deployed over $\Omega$. A vehicle $n$ follows the following mobility model [24]:
@@
To construct the estimate $\hat{f}$ of the unknown function $f$ over space and time, we want to collect measurements of the function $f$ with the AUVs. Given that the AUVs move according to (3), the measurements that we can use for the estimation are constrained in space and time by their $N$ trajectories. To formally state our problem formulation, we need to introduce some notation. Let us first define a vector of discrete time slots $\sigma=[0,\Delta,2\Delta,...,T\Delta]$ which we index with $\mathcal{K}=\{0,1,...,T\}$. In the rest of the paper, we are interested in updating the estimate $\hat{f}$ over the slots $\sigma[1],...,\sigma[T]$ with a slot granularity of $\Delta=30$ minutes.
@@
where $\epsilon_{x,t}$ is a generic noise term. Given space-time coordinates $A$, let $f(A)=\{f(x,t)+\epsilon_{x,t}:(x,t)\in A\}$ be the corresponding measurements, and let $\mathcal{M}_{k}^{N}=\{D_{k}^{N},f_{k}^{N}\}$.
@@ III Plume Monitoring with MARL
In this section, we describe the system architecture and the algorithm we design to solve problem (7). In Fig. 5 we provide a graphical illustration of the system architecture, and Algorithm 1 outlines the main control pipeline. The key rationales behind the system architecture are (i) offloading computation as much as possible from the AUVs to the server, prioritizing their endurance and (ii) relying only on intermittent and extremely low overhead communication. The pipeline at a given time slot $k$ is as follows: first, all agents dwell on the sea surface and establish a connection with the cloud server via Wi-Fi or GSM. During this communication, each AUV n=1,,Nn=1,...,N uplinks its latest measurements $\{Z_{k}^{n},f(Z_{k}^{n})\}$ collected over the past 30 minutes. In our implementation, the number of per-agent per-slot measurements $z_{k,n}$ depends on speed and ocean currents, averaging $5$ and limited to a maximum of $10$. Note that transmitting $5$ space-time coordinates and the corresponding measurements only involves transmitting $20$ float numbers, which - in the worst case - only requires $160$ bytes (always <500<500 bytes even with wireless transmission overhead). Upon receiving the new measurements $\{Z_{k}^{n},f(Z_{k}^{n})\}_{n=1}^{N}$, the server aggregates them with the previous set, resulting in $\mathcal{M}_{k}^{N}=\{D_{k}^{N},f_{k}^{N}\}=\mathcal{M}_{k-1}^{N}\cup\{Z_{k}^{(n)},f(Z_{k}^{(n)})\}_{n=1}^{N}$ and updates the salinity map estimate $\hat{f}(\cdot,\sigma[k]|\mathcal{M}_{k}^{N})$. Based on $\hat{f}$, the current agent trajectories $D_{k}^{N}$, and a vector containing the wind speed and direction $w_{k}$, the server determines the control policy $\pi(\hat{f},\mathcal{M}_{k}^{N},w_{k})$ for all AUVs n=1,,Nn=1,...,N for the next 30 minutes time slot, which results in direction $\{b_{k}^{(n)}\}_{n=1}^{N}$ and speed levels $\{v_{k}^{(n)}\}_{n=1}^{N}$. At that point, the server downlinks the commands to the AUVs, which resubmerge and resume sampling, starting to build $\{Z_{k+1}^{n},f(Z_{k+1}^{n})\}_{n=1}^{N}$. The estimator $\hat{f}(\cdot,\sigma[k]|\mathcal{M}_{k}^{N})$ and the policy $\pi(\hat{f},\mathcal{M}_{k}^{N},w_{k})$ form the core of our framework. We detail these two components in the following subsections, respectively.
@@ III-B Decision Making Module
To exploit the structure of the action space, we decompose $Q_{\theta}$ into two components (heads), which share a state-value function $V_{\bar{\theta}}(s)$ and are characterized by two different advantage functions $A_{\theta_{1},\text{dir}}(s,b)$, $A_{\theta_{2},\text{spd}}(s,v)$,
@@
where note that $\theta=\{\bar{\theta},\theta_{1},\theta_{2}\}$. For agent $n\in\{1,...,N\}$ in state $s_{k}^{(n)}$, action selection is then performed per head:
@@
which establishes direction and speed.
State construction. The state representation $s_{k}^{(n)}$ comprises three components. Following prior work on image-based policy inputs [30], we render both spatial estimates and trajectory traces into a fixed-resolution $3$-channels image, compressed by a convolutional neural network (CNN) into a compact embedding. Specifically, the three channels for the $n$-th agent are (i) a compressed image of the GP estimate of the salinity field $\hat{f}(\cdot,\sigma[k])$, (ii) one image marking in white over a black background recent measurement locations of the agent, $Z_{k}^{(n)},Z_{k-1}^{(n)},$... (iii) one image marking in white over a black background recent measurement locations for agent $n$ teammates, $\{Z_{k}^{(l)},Z_{k-1}^{(l)},...\}_{l\neq n}$. The trajectory marks intensities are log-scale weighted by recency. In addition, a wind vector $w_{k}=[\text{angle},\text{speed}]$, processed by a NN to produce an embedding, is concatenated with the rest of the network. See Fig. 5 for an illustration of the state construction.
@@ IV Simulation Study and Results
Dataset. We train our algorithm over a selection of $4$ months data from the year 2018 (February, April, October and December) and evaluate over different months from 2016-2018 (see Table I). Overall, we use a total of $\approx 6k$ time frames for the training, with a $30$ minutes time resolution. Each frame spans $\approx 5\times 10^{4}$ grid locations and measurement noise is i.i.d. $\mathcal{N}(0,\,0.01)$.
@@
Multi-agent performance. We now evaluate performance in the multi-agent setting, focusing on: (i) MSE gains provided by the multi-agent coordination when increasing $N$, and on the comparison with existing literature benchmarks; (ii) the energy efficiency improvements induced by varying the $\eta_{2}$ reward hyperparameter in (18) and the corresponding MSE-energy tradeoff. As a multi-agent benchmark, we implement the recent work in [32] that proposed an algorithm based on adaptive Voronoi partitioning of the space of interest to track a time-varying process. We empirically tune the exploration-exploitation hyperparameters in [32] and show the best results in terms of MSE. In Fig. 10, we compare the different schemes for $N=3$ and $N=6$ for March 2018. We first note how the adaptive Voronoi partition solution [32] significantly outperforms the strategic rotations solution. We also note that our algorithm provides maps with half of the MSE compared to [32], for both $N=3$ and $N=6$. In Fig. 12, we show a sequence of frames to illustrate the multi-agent control performance of the agents with respect to the salinity field $f$ evolution over a $15$ hour time interval. In Fig. 11 we show the endurance–MSE trade-off obtained by varying the reward weight $\eta_{2}$ in (18), using October 2017 and October 2018 as exemplars. Increasing $\eta_{2}$ biases the policy toward lower propulsion use, thus extending endurance at a modest cost in accuracy, while scaling from $N=3$ to $N=6$ both reduces MSE and more than doubles endurance. Table I summarizes results obtained with $\eta_{2}=50$, and further confirms that our algorithm generalizes across unseen (during training) years and seasonal regimes.